Find the center mass of the region bounded by y=e^2x and y=-cos(πx) between - 1/2 ≤x≤1/2.


1
Expert's answer
2021-04-26T07:24:28-0400

assuming uniform mass density so that "\u03c1(x,y)=k" for k a constant, the center of mass coincides with the centroid of the region;

we have

total mass; "m=\u03c1\u222b_{-1\/2}^{1\/2} [h(x)-g(x)]dx"

"=k\u222b_{-1\/2}^{1\/2}[e^2x-(-cos\u03c0x)dx]"

"=k[\\frac12 e^2x+\\frac1\u03c0 sin\u2061\u03c0x ]_{-1\/2}^{1\/2}"

"=[(\\frac12 e+\\frac 1\u03c0 sin\u2061 \\frac\u03c0x )-(\\frac 12 e^{-1}+ \\frac 1\u03c0 sin \\frac {-\u03c0}{2 } )]k"

"=(\frac 12 e-\frac12 e^{-1} +\frac 2π)k …(i)"

the moments of the region about the x- axis equal;

"m_x=\\frac k 2 \u222b_{-1\/2}^{1\/2}[(e^2x )^2+(-cos\u2061\u03c0x )^2] dx"

"=\\frac 12 k \u222b_{-1\/2}^{1\/2}(e^4\u03c0+cos^2\u2061\u03c0x)dx"

"=\\frac12k \u222b_{-1\/2}^{1\/2}[e^4x+\\frac 12 (1+cos\u20612\u03c0x )]dx"

"=\\frac12 k [\\frac14 e^2+\\frac 12 x+ \\frac14\u03c0 sin\u20612\u03c0x ]_{-1\/2}^{1\/2}"

"=\\frac 12 k [(\\frac14 e^2+\\frac14+\\frac14\u03c0 sin\u2061\u03c0 )-(\\frac 14 e^{-2}-\\frac14+\\frac14\u03c0 sin\u2061(-\u03c0))]"

"=\frac12 k [\frac14 e^2-\frac14 e^{-2}+\frac12]=k[\frac {e^4-1+2e^2 }{4e^2 }]…(ii)"

and moments of the region about y-axis ;

"=k\u222b_{-1\/2}^{1\/2} (xe^2x+xcos \u03c0x)dx"

"=k[\\frac12 xe^2x-\\frac14 e^2x+\\frac1\u03c0 x sin\u2061\u03c0x+\\frac1{\u03c0^2} cos\u2061\u03c0x ]_{-1\/2}^{1\/2}"

"=k(\frac12 e^{-1} )=\frac12 ke^{-1}"

the center of mass, ( "\bar x" , "\bar y" ) is the point

"\bar x=m_y/m" and "\bar y=m_x/m"

"\bar x=\frac {(\frac12 e^{-1})}{(k(\frac12 e-1/2 e^{-1}+\frac2π)}=\fracπ{πe^2-π+4 e}=0.102"

"\bar y =\frac{\frac12 k [\frac14 e^2-\frac14 e^{-2}+\frac12]} {k[\frac 12 e-- \frac 12 e^{-1}+\frac 2 \pi ))}= \frac {(πe^4-π+2πe^2)} {(4πe^3-4πe+16 e^2 )}=0.638"

"(0.102,0.638)" is the center of mass of the region